# The evolution of galaxy morphology from redshift $z=6$ to $3$: Mock JWST observations of galaxies in the ASTRID simulation

### v2

@@ 2.3.2 Image Creation
We used a constant approximation of the background radiance of $0.25$ $\rm MJy/sr$ across all filters, which is added to all observations. After the background is added, we simulate the photon noise of the observation using uncorrelated Poisson noise on the detection counts in each pixel. We calculate the expected detection counts in each pixel ($C_{pix,exp}$) using the following formula,
$C_{\rm pix,exp}=\frac{L_{\rm pix}}{{\rm sens}_{\rm avg}}*t,$ (2)
where $\rm sens_{\rm avg}$ is the average pixel sensitivity in $\rm(MJy/sr)/(DN/s)$, $L_{\rm pix}$ is the radiance in a given pixel in $\rm MJy/sr$ and $t$ is the exposure time in seconds. Using these expected detection counts in each pixel we perform a single Poisson sample for each pixel to determine its detected count, then convert back to $\rm MJy/sr$. The $\rm sens_{\rm avg}$ value for each filter is determined by averaging the "PHOTMJSR" values for each filter found in the JWST calibration and zeropoint table as of October 2022 (Gordon et al., 2022) [^2] Table is available at https://jwst-docs.stsci.edu/files/182256933/182256934/1/1669487685625/NRC_ZPs_0995pmap.txt.. The resulting $\rm sens_{\rm avg}$ values for the F [^4] W, F356W, F277W, F200W, and F150W filters are $0.383$, $0.391$, $0.453$, $1.914$, and $2.255$ $\rm MJy/sr$ respectively. The distinct jump between the F277W and F200W filters is due to the smaller pixel size of the short wavelength filters compared to the long wavelength filters.
@@ 3 Results
We use the above imaging pipeline to create synthetic images of $\sim 355,000$ subhalos from redshift 3 to 6 with five NIRCAM filters (F150W, F200W, F277W, F356W, F [^4] W). We measure a number of morphological and size properties of these galaxies in order to compare this dataset to existing observations, and similar simulation based data sets, and discern as much as possible about the morphologies of galaxies in this redshift range.
@@ 3.1.1 Sersic Fitting
The distribution of Sersic indices for our galaxy sample is shown in Figure 5. We can see that the overall distribution remains largely unchanged across the redshift range in our sample. Traditionally, Sersic indices near $n=1$ are associated with disk galaxies, and larger Sersic indices are associated with elliptical galaxies ($n=4$ produces the de Vaucouleurs profile). The distribution of Sersic indices for our entire sample has a mean of $n=0.75$ and $\sigma=0.21$, with only $0.12\%$ of galaxies having a Sersic index of 2 or higher. Observational datasets for galaxies at these redshifts indicate that disk galaxies should have a narrow distribution centered near $n=1.0$ while elliptical galaxies should have a broader distribution centered between $n=2.0$ and $n=3.0$ (Kartaltepe et al., 2023; Vika et al., 2015; Kartaltepe et al., 2015). This indicates we may be capturing elliptical galaxies that fall on the lower end of the broad distribution, but are likely missing many of the higher Sersic index galaxies. That said, our sample matches quite well with other samples made from simulations. In the first column of figure 7 we can see our distribution of Sersic indices compared with other samples including those made form Illustris TNG50 and TNG100 (Costantin et al., 2023; Rose et al., 2023), and our distribution is comparable to theirs, while being slightly narrower.
@@ 3.1.3 Comparison with other datasets
[…]
@@ 3.2.2 Observation Drizzling
We also examine the impact that drizzling and exposure time have on these distributions by fitting single exposures with exposure time $t=945s$ compared to the drizzled images made from 3 exposures. For the subsample we examine this has a negligible effect on the mean of the distribution for both Sersic index ($n=0.74\pm 0.28,n=0.71\pm 0.23$ for undrizzled and drizzled images respectively) and the Sersic effective radius ($r_{e}=1.91\pm 0.77kpc,r_{e}=1.95\pm 0.74kpc$ respectively). One the other hand, the undrizzled images are significantly more likely to have issues during the fitting process. 98% of the drizzled observations are able to be fit without issue, while only 90% of the undrizzled observations of the same galaxies are properly fit. All of the galaxies that are well fit using the drizzled observations, but are unable to be fit using the undrizzled observations have a mass of $log{M_{*}/\,M_{\odot}}=9.7$ or less. This shows the value of the increased sensitivity of JWST’s instruments, and longer exposure drizzled observations for capturing lower mass dim galaxies.
@@ 4.1.1 Morphology-Mass and Size-Mass Relations
and the results can be found in table 2. There is a significant change in the relationship between $n$ and mass as time progresses. The relationship is consistently positive, and generally increases as redshift decreases ($\alpha_{n}=0.100,0.066,0.108,$ and $0.172$ for $z=6,5,4,$ and $3$ respectively). These relations are the most drastic at the high mass end, as the relations are largely flat for galaxies with $log(M_{*}/\,M_{\odot})\leq 10.0$, and the high number of galaxies at the low mass end tends to dominate the fit. The high mass behavior exhibits a consistent trend of galaxies in the same mass bin having higher Sersic indices at lower redshift, indicating that the morphology of high mass halos evolves over this redshift range. We perform the same linear fit outlined above on the population of galaxies with $log(M_{*}/\,M_{\odot})\geq 10.25$ to quantify this evolution. This can also be seen in table 2. Specifically this indicates that the population of high mass galaxies contains more disks and irregular galaxies at high redshift, and that the population shifts toward more elliptical galaxies as time progresses.
There is a great deal of work using observations (Whitney et al., 2021; Bluck et al., 2012; Tamburri et al., 2014), simulations (Rodriguez-Gomez et al., 2015), and semi-analytical models (Hopkins et al., 2010) that show galaxy merger rates increase with redshift which would lead to more high mass irregular galaxies with low Sersic indices at higher redshifts. There is also work that shows this same trend of high mass galaxies having lower Sersic indices and being more disk-like at higher redshifts (Buitrago et al., 2013; Ormerod et al., 2023; Shibuya et al., 2015). Both of these effects would lead to lower average Sersic indices for high mass galaxies at higher redshifts, so the presence of that trend in our data supports them but this analysis alone is not enough to differentiate between the two.
@@
Another noticeable feature of the $r_{e}$ vs mass plot is that all 4 redshifts have a fairly flat relationship between mass and radius for galaxies below $\log(M_{*}/\,M_{\odot})=10.5$, and display an upturn after that point. This upturn is barely noticeable for $z=6$ and $z=5$ but becomes more evident for $z=4$ and $z=3$. Similar upturns at the high mass end are seen in observations and supported by theory regarding the growth mechanisms for galaxies at different masses (Nedkova et al., 2021; Mosleh et al., 2020; Shen et al., 2003; Mosleh et al., 2013; Mowla et al., 2019a).
@@ 4.2.2 Comparing CAS and Gini-$M_{20}$ measurements to Sersic Fits
The Gini-$M_{20}$ plane colored by Sersic index has a trend of increasing Sersic index up and to the right which is roughly perpendicular with the boundary between disks and ellipticals, indicating that they are very correlated. This shows that classifications based on Sersic index, and based on the Gini-$M_{20}$ plot should have similar results. The galaxies that fall in the elliptical portion of the Gini-$M_{20}$ plot have significantly higher Sersic indices ($n=1.3050\pm 0.0068$) than those in the disk region ($n=0.74692\pm 0.00044$). The low number of galaxies classified as elliptical based on the Gini-$M_{20}$ boundaries does align with the low number of galaxies with high Sersic indices in our sample.
@@
Mergers -1.2668 0.5174 0.0064 0.0021
Disks -1.5316 0.4567 0.0005 0.0002
Ellip/S0 -1.6030 0.4489 0.0003 0.0001
@@
Mergers 2.7704 0.1958 0.0069 0.0024
Disks 2.4679 0.0606 0.0004 0.0001
Ellip 2.9737 0.0773 0.0046 0.0010

### v3

@@ 2.3.2 Image Creation
We used a constant approximation of the background radiance of $0.25$ $\rm MJy/sr$ across all filters, which is added to all observations. After the background is added, we simulate the photon noise of the observation using uncorrelated Poisson noise on the detection counts in each pixel. We calculate the expected electron counts in each pixel ($N_{pix,exp}$) using the following formula,
$N_{\rm pix,exp}=\frac{L_{\rm pix}}{{\rm sens}_{\rm avg}}*\rm e_{\rm gain}*t,$ (2)
where $\rm sens_{\rm avg}$ is the average pixel sensitivity in $\rm(MJy/sr)/(DN/s)$, $L_{\rm pix}$ is the radiance in a given pixel in $\rm MJy/sr$, $\rm e_{\rm gain}$ is the gain between electrons and digital numbers in $\rm e^{-}/DN$ and $t$ is the exposure time in seconds. Using these expected electron counts in each pixel we perform a single Poisson sample for each pixel to determine its electron count, then convert back to $\rm MJy/sr$. We used a value of 2.0 for $\rm e_{\rm gain}$ for all filters. The $\rm sens_{\rm avg}$ value for each filter is determined by averaging the "PHOTMJSR" values for each filter found in the JWST calibration and zeropoint table as of October 2022 (Gordon et al., 2022) [^2] Table is available at https://jwst-docs.stsci.edu/files/182256933/182256934/1/1669487685625/NRC_ZPs_0995pmap.txt.. The resulting $\rm sens_{\rm avg}$ values for the F [^4] W, F356W, F277W, F200W, and F150W filters are $0.383$, $0.391$, $0.453$, $1.914$, and $2.255$ $\rm(MJy/sr)/(DN/s)$ respectively. The distinct jump between the F277W and F200W filters is due to the smaller pixel size of the short wavelength filters compared to the long wavelength filters.
@@ 3 Results
We use the above imaging pipeline to create synthetic images of $\sim 350,000$ subhalos from redshift 3 to 6 with five NIRCAM filters (F150W, F200W, F277W, F356W, F [^4] W). We measure a number of morphological and size properties of these galaxies in order to compare this dataset to existing observations, and similar simulation based data sets, and discern as much as possible about the morphologies of galaxies in this redshift range.
@@ 3.1.1 Sersic Fitting
The distribution of Sersic indices for our galaxy sample is shown in Figure 5. We can see that the overall distribution remains largely unchanged across the redshift range in our sample. Traditionally, Sersic indices near $n=1$ are associated with disk galaxies, and larger Sersic indices are associated with elliptical galaxies ($n=4$ produces the de Vaucouleurs profile). The distribution of Sersic indices for our entire sample has a mean of $n=0.72$ and $\sigma=0.21$, with only $0.10\%$ of galaxies having a Sersic index of 2 or higher. Observational datasets for galaxies at these redshifts indicate that disk galaxies should have a narrow distribution centered near $n=1.0$ while elliptical galaxies should have a broader distribution centered between $n=2.0$ and $n=3.0$ (Kartaltepe et al., 2023; Vika et al., 2015; Kartaltepe et al., 2015). This indicates we may be capturing elliptical galaxies that fall on the lower end of the broad distribution, but are likely missing many of the higher Sersic index galaxies. That said, our sample matches quite well with other samples made from simulations. In the first column of figure 7 we can see our distribution of Sersic indices compared with other samples including those made form Illustris TNG50 and TNG100 (Costantin et al., 2023; Rose et al., 2023), and our distribution is comparable to theirs, while being slightly narrower.
@@ 3.1.3 Comparison with other datasets
When comparing to these datasets we check the impact that the galaxy stellar mass function (GSMF) and redshift distribution has on the result, as our population has a slightly different GSMF than the CEERS dataset itself, as seen in figure 2, and our population is confined to integer redshifts rather than having a continuous redshift distribution. We weight our dataset by galaxy mass and redshift in order to match the CEERS sample as closely as possible in those two metrics and compare with our results shown in figure 7. We see minimal shifts in all metrics. For example the mean Sersic index in the redshift 3 to 4 band shifts by 0.01 after weighting.
@@ 3.2.2 Observation Drizzling
We also examine the impact that drizzling and exposure time have on these distributions by fitting single exposures with exposure time $t=945s$ compared to the drizzled images made from 3 exposures. For the subsample we examine this has a negligible effect on the mean of the distribution for both Sersic index ($n=0.74\pm 0.27,n=0.71\pm 0.22$ for undrizzled and drizzled images respectively) and the Sersic effective radius ($r_{e}=1.95\pm 0.79\rm kpc,r_{e}=1.95\pm 0.74\rm kpc$ respectively). On the other hand, the undrizzled images are slightly more likely to have issues during the fitting process. 99% of the drizzled observations are able to be fit without issue, while only 96% of the undrizzled observations of the same galaxies are properly fit. All of the galaxies that are well fit using the drizzled observations, but are unable to be fit using the undrizzled observations have a mass of $\rm log{M_{*}/\,M_{\odot}}=9.7$ or less. This shows the value of the increased sensitivity of JWST’s instruments, and longer exposure drizzled observations for capturing lower mass dim galaxies.
@@ 4.1.1 Morphology-Mass and Size-Mass Relations
and the results can be found in table 2. There is a significant change in the relationship between $n$ and mass as time progresses. The relationship is consistently positive, and generally increases as redshift decreases ($\alpha_{n}=0.112,0.079,0.097,$ and $0.142$ for $z=6,5,4,$ and $3$ respectively). These relations are the most drastic at the high mass end, as the relations are largely flat for galaxies with $log(M_{*}/\,M_{\odot})\leq 10.0$, and the high number of galaxies at the low mass end tends to dominate the fit. The high mass behavior exhibits a consistent trend of galaxies in the same mass bin having higher Sersic indices at lower redshift, indicating that the morphology of high mass halos evolves over this redshift range. We perform the same linear fit outlined above on the population of galaxies with $log(M_{*}/\,M_{\odot})\geq 10.25$ to quantify this evolution. This can also be seen in table 2. Specifically this indicates that the population of high mass galaxies contains more disks and irregular galaxies at high redshift, and that the population shifts toward more elliptical galaxies as time progresses.
There is a great deal of work using observations (Whitney et al., 2021; Bluck et al., 2012; Tamburri et al., 2014), simulations (Rodriguez-Gomez et al., 2015), and semi-empirical models (Hopkins et al., 2010) that show galaxy merger rates increase with redshift which would lead to more high mass irregular galaxies with low Sersic indices at higher redshifts. There is also work that shows this same trend of high mass galaxies having lower Sersic indices and being more disk-like at higher redshifts (Buitrago et al., 2013; Ormerod et al., 2023; Shibuya et al., 2015). Both of these effects would lead to lower average Sersic indices for high mass galaxies at higher redshifts, so the presence of that trend in our data supports them but this analysis alone is not enough to differentiate between the two.
@@
Another noticeable feature of the $r_{e}$ vs mass plot is that all 4 redshifts have a fairly flat relationship between mass and radius for galaxies below $\log(M_{*}/\,M_{\odot})=10.5$, and display an upturn after that point. This upturn is noticeable for $z=4-6$ but becomes most pronounced for $z=4$. Similar upturns at the high mass end are seen in observations and supported by theory regarding the growth mechanisms for galaxies at different masses (Nedkova et al., 2021; Mosleh et al., 2020; Shen et al., 2003; Mosleh et al., 2013; Mowla et al., 2019a).
@@ 4.2.2 Comparing CAS and Gini-$M_{20}$ measurements to Sersic Fits
The Gini-$M_{20}$ plane colored by Sersic index has a trend of increasing Sersic index up and to the right which is roughly perpendicular with the boundary between disks and ellipticals, indicating that they are very correlated. This shows that classifications based on Sersic index, and based on the Gini-$M_{20}$ plot should have similar results. The galaxies that fall in the elliptical portion of the Gini-$M_{20}$ plot have significantly higher Sersic indices ($n=1.2924\pm 0.0097$) than those in the disk region ($n=0.71072\pm 0.00040$). The low number of galaxies classified as elliptical based on the Gini-$M_{20}$ boundaries does align with the low number of galaxies with high Sersic indices in our sample.
@@
Mergers -1.2500 0.5163 0.0053 0.0018
Disks -1.5148 0.4520 0.0005 0.0001
Ellip/S0 -1.5927 0.4544 0.0003 0.0001
@@
Mergers 2.7347 0.2143 0.0064 0.0022
Disks 2.4441 0.0669 0.0004 0.0001
Ellip 2.9794 0.0867 0.0053 0.0012